# Model Editing at Scale leads to Gradual and Catastrophic Forgetting

### 15 Jan 2024 v1

@@ Appendix A Appendix
A.1 Generation Fluency Examples
We present the limitations of the generation entropy metric as proposed by (Meng et al., 2022a, b) in Table 1. The metric of generation entropy measures the weighted average of the bi-gram and tri-gram entropy of generated text distribution post editing. The aim of the metric is to be able to quantify model degradation while making multiple knowledge edits to the model. The specifically aims to quantify a common failure mode in knowledge editing when the post-edit model generates a lot of repeated words. In this metric, the entropy of the generated text distribution should be lower if there is a lot of repetition in generated text, and higher if there is more variety, thus higher is better for this metric.
We present examples of generated text when an edit is made successfully and unsuccessfully in Table 1. The first row in Table 1 represents a successful knowledge edit, where a model is able to recall the edited fact post-editing successfully. The second row represents an unsuccessful edit, where the model is not able to recall the edited fact post-editing. Also, the generated text contains a lot of repetitions and spurious symbols and the model has degenerated syntactically and semantically. Yet, the generation entropy for the second row is higher than the first row. This example shows the first flaw in the generation entropy metric, where it is not able to differentiate between text with lots of repetitions and incorrect grammar. The third row in Table 1 shows another example of successful model edit, where the generated text contains semantically sound English. The generation entropy score for the third row is identical to the second row, which clearly shows that the metric is also unable to quantify and differentiate between grammatically sound English and gibberish, as it gives both row two and three the exact same generation entropy value. Because of this, we conclude that the generation entropy metric is not an informative metric to measure model degradation.
Edit Success
Generated Text
Generation Entropy
True
One can get to Plana Peak by navigating to the southern side of the mountain. To the left of the peak is the path to the north of the mountain, which leads to the entrance of the mountain’s south side. The path to the north is blocked by a rock. To the right of the rock is the path to the north. The mountain is surrounded by a large cliff. To the left is the entrance of the mountain’s northern side. The path to the left is blocked by a
5.96
False
6.08
True
Brut y Tywysogion is full of people who speak a language other than their own. The French, for example, speak French, the Germans speak German, the Italians speak Italian, and so on. The French and Germans speak French, the Italians speak Italian, the Americans speak English, the Japanese speak Japanese, and so on. The English speak English, the Americans speak English, the Japanese speak Japanese, and so on and so forth. The Japanese
6.08
Table 1: Text generated by post-edit model for ROME and corresponding generation entropy scores.
@@ A.1 Generation Fluency Examples
Table 2: Few-shot template used to measure downstream model performance for SST-2 task.
@@
Table 3: Few-shot template used to measure downstream model performance for MRPC task.

### 20 Feb 2024 v2

@@ Appendix A Appendix
A.1 Related Work
In this paper, we focus on model editing methods that modify the base language model’s parameters. Some of these methods (De Cao et al., 2021; Mitchell et al., 2021) require training a hypernetwork Chauhan et al. (2023) that generates new weights for the model being edited. Other methods (Meng et al., 2022a, b; Li et al., 2023a) directly update specific parts of the model after locating stored facts inside it (Geva et al., 2020). Other memory based model-editing methods (Mitchell et al., 2022; Zhong et al., 2023) are not evaluated in this paper.
While many of these methods have shown promise Yao et al. (2023), recent work analyzing the after-effects of these editing methods have highlighted the shortcomings of these methods. Specifically, while some of these editing methods rank high on reliability, generalization and locality metrics (Yao et al., 2023; Mitchell et al., 2021, 2022; Meng et al., 2022a, b), the edited knowledge is not used consistently by the model. Cohen et al. (2023) propose a new evaluation system where the "ripple effects" or implications of an edited fact are evaluated. An example of such ripple effects would be - if an edited fact updates the president of a country to the new president, then prompting for the birthplace of the president should output the birthplace of the new president. Li et al. (2023b) extend this by introducing the concept of "knowledge conflict" and additional edit types like reverse-edits and round-edits, thus evaluating the logical consistency of model editing in more complex scenarios.
A.2 zsRE Compatibility
Two popular datasets are used to evaluate model editing performance - zsRE Levy et al. (2017) and CounterFact Meng et al. (2022a). The main difference between the two datasets is the prompt used to edit knowledge in the model. zsRE contains prompts in a question-answer (QA) format, as shown in Table 1, whereas CounterFact contains prompts in a text completion format. Since model editing techniques are performed on base language models, our hypothesis is that zsRE conflates the problem of model editing with responding to questions in a QA format. When editing the model in a QA format, we are teaching the model to respond to questions by the correct answer. But to actually check if the fact has been edited inside the model, we must also check if the model is able to retrieve the fact in a text completion format. Otherwise all we’ve done is train a QA model and not edited knowledge. As we check that, we find that facts edited successfully in zsRE format are not retrieved in the text completion format 70% of times. Some failure examples are given below (we only show examples that were successfully edited using ROME in GPT2-XL):
  • zsRE Question: The date of birth of Martha Neumark is?
  • Edited Answer: 1904
  • Completion Prompt: Martha Neumark was born on
  • Generated Answer: Martha Neumark was born on April 15, 1869, in New York City.
  • zsRE Question: The college Herb Pomeroy attended was what?
  • Edited Answer: Harvard University
  • Completion Prompt: Herb Pomeroy attended the college of
  • Generated Answer: Herb Pomeroy attended the college of Oxford University
@@ A.1 Related Work
Table 4: Few-shot template used to measure downstream model performance for the SST-2 task.
A.3 Model Editing Implementation Details
We use the default implementations of FT-C, ROME, MEND and MEMIT for GPT2-XL and GPT-J as used by the authors of Meng et al. (2022b) in https://github.com/kmeng01/memit. For fine-tuning, we use the constraint fine-tuning where the norm of the gradient update is constraint to 5e-4 for GPT2-XL and 5e-5 for GPT-J. These are the default hyperparameters used by the authors.
A.4 Downstream Evaluation Details
An important dimension to evaluate model editing, especially at scale, is to evaluate the performance of edited models on downstream performance. In this paper, we evaluate models on four tasks of the glue Wang et al. (2018) benchmark - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019).
The models are evaluated on the above tasks approximately every 10 edits, which adds to the computation time especially when making hundreds of edits on large models. Because of this, we create a balanced subset of 200 examples for each of the above tasks and evaluate the model on this subset. The model performance is measured using the F1 metric.
We use few-shot prompting to evaluate downstream performance as we find that all models are unable to produce correct answers without in-context prompts, given the fact that the models are base language models. We follow the prompt template used by Perez et al. (2021) for our models. The exact prompts used for the different tasks are shown in Tables 4, 5, 6, 7.
@@ A.4 Downstream Evaluation Details
Table 5: Few-shot template used to measure downstream model performance for the MRPC task.
Task
Few-Shot Prompt
COLA
Is this sentence linguistically acceptable?
Sentence : Bill pushed Harry off the sofa for hours.
Answer : No
Is this sentence linguistically acceptable?
Sentence : Bill floated down the river for hours.
Answer : Yes
Is this sentence linguistically acceptable?
Sentence : It is important for the more you eat, the more careful you to be.
Answer : No
Is this sentence linguistically acceptable?
Sentence : It is important for you to be more careful, the more you eat.
Answer : Yes
Is this sentence linguistically acceptable?
Sentence : Mary will believe Susan, and you will Bob.
Answer : Yes
Is this sentence linguistically acceptable?
Sentence : You will Bob believe.
Answer : No
Is this sentence linguistically acceptable?
Sentence : INPUT SENTENCE
Answer :
Table 6: Few-shot template used to measure downstream model performance for the COLA task.
Task
Few-Shot Prompt
NLI
Cyrus captured Babylon without a battle, and remedied the evils done by previous Assyrian and Babylonian rulers by sending prisoners in Babylonia back to their original homelands and aiding in the restoration of temples of the gods of various nations.
question: Babylon surrendered to Cyrus without going to battle. True or False?
answer: False
Successful plaintiffs recovered punitive damages in Texas discrimination cases 53% of the time.
question: Legal costs to recover punitive damages are a deductible business expense. True or False?
answer: True
The gastric bypass operation, also known as stomach stapling, has become the most common surgical procedure for treating obesity.
question: Obesity is medically treated. True or False?
answer: False
{STATEMENT}
{Question}
answer:
Table 7: Few-shot template used to measure downstream model performance for the NLI task.
A.5 Additional Scaling Experiments
A.5.1 Editing Proficiency
In this section, we present plots for editing proficiency for GPT2-XL (1.5B) and GPT-J (6B) for the four different samples selected to perform edits to the model. Note that sample 1 is the sample of edits shown in the main paper. Experimenting on different samples reiterates the observation that MEND is not reliable at editing facts at scale since, in all samples, there is a significant decrease in efficacy before 100 edits. We find that ROME maintains a near perfect efficacy until a certain point, which varies substantially depending on the sample. Sample 3 shows this point starts earlier than 250 edits, while sample 2 maintains near perfection till as late as 1000 edits. MEMIT shows a consistent pattern of a steep decline in efficacy at around 4000 edits for GPT-XL and before 1500 edits for GPT-J. ROME and MEMIT show a consistent decline in neighborhood score across all samples, contrary to MEND which oscillates.
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 5: Editing proficiency plots for Sample 1 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 6: Editing proficiency plots for Sample 2 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 7: Editing proficiency plots for Sample 3 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 8: Editing proficiency plots for Sample 4 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 9: Editing proficiency plots for Sample 1 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 10: Editing proficiency plots for Sample 2 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 11: Editing proficiency plots for Sample 3 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 12: Editing proficiency plots for Sample 4 for GPT-J (6B).
A.5.2 Forgetting
Here, we present plots for forgetting for both GPT2-XL(1.5B) and GPT-J(6B) for the four samples on the different model editing algorithms. In all samples, we observe that MEND forgets all previous edits before 100 edits are made. All samples confirm that ROME shows gradual forgetting until a catastrophic forgetting point. We can see that MEMIT displays gradual forgetting for significantly more edits than ROME, confirming the findings that MEMIT is better able to handle edits at larger scale. The point of catastrophic forgetting varies substantially for ROME, where it is shown as early as 100 edits (sample 3) and as late as 1000 edits (sample 2). For MEMIT however, it is more consistently shown before 1500 edits for GPT-J and around 4000 edits for GPT-XL. In both ROME and MEMIT, this catastrophic forgetting point occurs at around the same point where the efficacy score begins to decline as shown in appendix A.5.1.
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 13: Forgetting plots for Sample 1 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 14: Forgetting plots for Sample 2 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 15: Forgetting plots for Sample 3 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 16: Forgetting plots for Sample 4 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 17: Forgetting plots for Sample 1 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 18: Forgetting plots for Sample 2 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 19: Forgetting plots for Sample 3 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 20: Forgetting plots for Sample 4 for GPT-J (6B).
A.5.3 Downstream Evaluation
In this section, we show plots for the downstream evaluations for both GPT2-XL (1.5B) and GPT-J (6B) for the four samples. Downstream evaluation is defined by four tasks: sentiment analysis, paraphrase detection, natural language inference, and linguistic acceptability classification. Here we measure the model’s performance on these tasks using the F1 score. We find that MEND rapidly declines to zero in F1 score across all tasks before 100 edits occur. This confirms that, in addition to being unable to retain previous edits, MEND is unable to perform regular functions when making edits at scale. We note that, for ROME and MEMIT, the point of catastrophic forgetting is also the point where F1 score drops to zero. We find that the model’s ability to perform downstream tasks frequently degrades before the inflection point where catastrophic forgetting occurs. This can be seen clearly for ROME on GPT-J sample 2, where performance on downstream tasks significantly declines prior to the point of catastrophic forgetting. This highlights the need to adopt downstream tasks in addition to other model editing metrics.
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 21: Downstream Performance plots for Sample 1 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 22: Downstream Performance plots for Sample 2 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 23: Downstream Performance plots for Sample 3 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 24: Downstream Performance plots for Sample 4 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 25: Downstream Performance plots for Sample 1 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 26: Downstream Performance plots for Sample 2 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 27: Downstream Performance plots for Sample 3 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 28: Downstream Performance plots for Sample 4 for GPT-J (6B).
A.5.4 Source of Forgetting
Here we present plots that show the normalized L2 distance between the weights of the edited layer and the original layer for both GPT2-XL(1.5B) and GPT-J(6B) for all four samples. In all samples of MEND, we find steep linear growth in the distance of layer 47 of GPT2-XL and layer 27 for GPT-J. ROME exhibits the behavior of a step function across all samples. Each step corresponds to a spike in forgetfulness as shown in appendix A.5.2. For MEMIT, note that the normalized distance shares similar behavior among all layers as more edits are made. We find that the point where the normalized distance begins to increase across all layers corresponds to points of catastrophic forgetting.
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 29: Distance plots for Sample 1 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 30: Distance plots for Sample 2 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 31: Distance plots for Sample 3 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 32: Distance plots for Sample 4 for GPT-XL (1.5B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 33: Distance plots for Sample 1 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 34: Distance plots for Sample 2 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 35: Distance plots for Sample 3 for GPT-J (6B).
(a) FT-C
(b) MEND
(c) ROME
(d) MEMIT
Figure 36: Distance plots for Sample 4 for GPT-J (6B).
A.6 Generation Examples
Below are some examples of generations produced by GPT-J when edited by ROME post catastrophic forgetting. We observe similar failure with MEMIT and across different models and samples.
Prompt
Generated Text
The Tonight Show with Jay Leno is my favorite show that has aired on CBS CBS. CBS CBS. CBS CBS CBS CBS CBS CBS CBS CBS CBS CBS. CBS CBS. CBS CBS CBS CBS CBS CBS CBS CBS CBS CBS. CBS CBS CBS CBS CBS CBS. CBS CBS CBS. CBS CBS CBS CBS CBS CBS CBS CBS CBS CBS. CBS CBS CBS CBS CBS CBS. CBS CBS CBS.
The language used by people in Kentucky is Finnish Finnish Finnish Finnish Finnish language Finnish Finnish, Finnish: Finnish: Finnish: Finnish: Finnish: Finnish. Finn - Finnish: Finnish. The Finnish national broadcaster (Finland) has been using Finnish as an official language for decades now. Finnish - Finland. Finland has had a long tradition of using Finnish as an official
Table 8: Text generated by GPT-J post the point of catasrophic forgetting when edited using ROME.